Ticket Number: HEA-NOC/20070928-12 Ticket Status: OPEN
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20070928 15:43 UTC+1 Problem Start: 20071003 17:00 UTC+1
Ticket Update: 20070928 15:43 UTC+1 Problem End:
Site/line: UCD, BGP Routing
Migrating UCD from OSPF to BGP
20070928 15:29 OMG Sent BGP configs for IPv4 (unicast & multicast) and
IPv6 (unicast) to NOR for bringing up the link. BGP peerings to be
established next Tuesday, but OSPF will remain in place as primary routing
protocol for 1 week.
This ticket can be monitored at http://www.hea.net/tickets/20070928-12
HEAnet Limited. Registered in Ireland, No. 275301.